OpenStudy (zzr0ck3r):

Abstract Algebra H is a subset of R x R H := {(x,y) s.t. y = 2x} Prove H is a normal subgroup of R x R.

OpenStudy (zzr0ck3r):

I don't understand what the operation is and if its multiplication if (x,y) is in H and (x',y') is in H the y = 2x and y' = 2x' so yy' = 2*2xx' and thus not in H....

OpenStudy (kinggeorge):

Assuming R is the real numbers, the operation must be addition. Otherwise things go very badly with 0.

OpenStudy (zzr0ck3r):

right good call.

OpenStudy (zzr0ck3r):

so it would be the ordered pair (x+x',y+y') and we would have y+y' = 2(x+x')?

OpenStudy (kinggeorge):

yes.

OpenStudy (zzr0ck3r):

and since R is commutative with addition H is a subgroup of an Abelian group and thus H is normal?

OpenStudy (zzr0ck3r):

every subgroup of an abelian group is normal.

OpenStudy (kinggeorge):

That would work.

OpenStudy (zzr0ck3r):

word. ty again kind sir:)

OpenStudy (kinggeorge):

You're welcome. Also, you still need to prove that it's a subgroup, but I assume you can handle that/already have.

OpenStudy (zzr0ck3r):

yeah for sure.